OpenStudy (anonymous):

What is y = (x − 3)(4x + 2) written in standard form?

OpenStudy (ivettef365):

FOIL

OpenStudy (anonymous):

ok uh... hold on

OpenStudy (anonymous):

Then combine like-terms.

OpenStudy (anonymous):

y=4x^2 + 2x - 7x - 6 then y=4x^2 - 5x - 6 ?

hero (hero):

y = x(4x + 2) - 3(4x + 2) = 4x^2 + 2x - 12x - 6 = 4x^2 - 10x - 6

OpenStudy (anonymous):

10x?

OpenStudy (anonymous):

Yeah, I got 4x^2-10x-6. You must have multiplied two terms together wrong.

OpenStudy (anonymous):

huh... explain that to me please

OpenStudy (anonymous):

\[(x-3)(4x-12)=4x^2+2x-12x-6\]

hero (hero):

@brittanny_, I showed you the steps

OpenStudy (anonymous):

Combine \[2x-12x\] to get \[-10x\]

OpenStudy (anonymous):

ok that makes sense now. thanks!

hero (hero):

y = (x - 3)(4x + 2) = x(4x + 2) - 3(4x + 2) = 4x^2 + 2x - 12x - 6 = 4x^2 - 10x - 6

hero (hero):

@britanny_ analyze that for a minute until you understand it.

hero (hero):

It's just the distributive property, which is the same thing as FOIL

OpenStudy (anonymous):

You must have added \[-3\] and \[4x\] instead of multiplying them together to get \[12x\]

OpenStudy (anonymous):

okay that makes sense. thank you!

OpenStudy (anonymous):

ah yeah i did do that...

OpenStudy (anonymous):

urg! i wanna give both of you medals!